The product of reeta's age(in years) five years ago and her age nine years later is 15, the present age of reeta is?
Answer by Jerome952(2)

If Reetah's age now is X
Reetah's age five years ago = X-5
Reetah's age in 9years to come=X +9
the product of Reetah's age in 5years before and 9years to come
(X-5)(X +9)=15
X^2 +9x-5X-45=15
collecting like terms to form a proper quadratic equation
X^2 +4X-60=0
solving the quadratic equation
X^2 +10X-6X-60=0
X(X +10)-6(X +10)=0
(X-6)(X +10)=0
therefor X=6 or X=-10
but we shall choose a positive answer which is X=6years
So Reetah's age now is 6years